OpenStudy (anonymous):

1/4x - 1/8x = 3 - 1/16x

OpenStudy (anonymous):

16

OpenStudy (anonymous):

\[\frac{1}{4} - \frac {1}{8} = \frac{1}{8} + \frac{1}{16}=\frac{3}{16} \]

OpenStudy (anonymous):

\[\frac{1}{3} \times \frac{3}{16}=16\]

OpenStudy (anonymous):

First group like terms on one side, 1/4x - 1/8x + 1/16x = 3 Then find the l.c.m of the L.H.S l.c.m = 16x ( 4 - 2 + 1)/ 16x = 3 3/16x = 3 Multiply both sides by 16x to do away with the denominator 3 = 16x * 3 3 = 48x Divide both sides by 48 to allow x to stand alone x = 3 / 48 x = 1 / 16 or 0.0625

OpenStudy (anonymous):

Thank you both

OpenStudy (anonymous):

Got another....

OpenStudy (anonymous):

When you fix the answer of x into the equation, it should be true. e.g 1/4(0.0625) - 1/8(0.0625) = 3 - 1/16(0.0625) the L.H.S should always be equal to the R.H.S

OpenStudy (anonymous):

@chuk32chuk , check your answer again. i think you might be wrong

OpenStudy (anonymous):

.25-.125=.125+.0625 (the 1/16) =.1875 Then we have .1875x=3 3/.1875 = 16